Is there a sphere passing through (4,0,1),(10,-4,9),(-5,6,-11)and (1,2,3)?If so,find its equation.
1
Expert's answer
2020-06-02T19:15:16-0400

We remind that an equation of a sphere has a form:

"(x-a)^2+(y-b)^2+(z-c)^2=r^2,\quad a,b,c,r\in{\mathbb{R}}"

After substitution of points we get a system of four equations:

"\left\{ 
\begin{matrix}
 (4-a)^2+b^2+(1-c)^2=r^2 \\
 (10-a)^2+(4+b)^2+(9-c)^2=r^2 \\
 (5+a)^2+(6-b)^2+(11+c)^2=r^2 \\
 (1-a)^2+(2-b)^2+(3-c)^2=r^2 \\
\end{matrix}\right."

We subtract the first equation from the others, simplify expressions and receive:

"\left\{ 
\begin{matrix}
 12a-8b+16c=180 \\
 -18a+12b-20c=165 \\
 6a-4b-4c=3 \\
\end{matrix}\right."

We multiply the last equation by (-2) and add to the first one. Then we get "24c=174" . This implies "c=7.25". Now we multiply the last equation by 3 and add to the second. We receive "-32c=174". Then "c=-5.4375" . Thus, we received different values of "c". Therefore,

the system has no solutions. Thus, there is no sphere passing through points (4,0,1),(10,-4,9),(-5,6,-11)and (1,2,3).